4) Townville Bank offers a savings account that pays 3.25% annual interest compounded monthly. a) What...

  1. 4) Townville Bank offers a savings account that pays 3.25% annual interest compounded monthly.

    a) What is the yield (APY) on this account? Round to the nearest hundredth of a percent. [5 pt]

    b) If Amy makes a deposit of $15,000 into this account, how much will be in her account after 8 years?

    [8 pt]
Answer #1

ANSWER IN THE IMAGE ((YELLOW HIGHLIGHTED). FEEL FREE TO ASK ANY DOUBTS. THUMBS UP PLEASE.

normal method:

apy={ [1+(3.25%/12)]^(12)}-1= 3.30%

future value= 15000*(1.0329885)^8 = 19447.12

balance after 8 yrs= 19447.12

excel:

MQ416 fx X MP =MQ412 MQ MO MR MS MI MU 407 408 409 410 411 rate APY 3.25% 3.30% =EFFECT(MQ411,12) 412 413 414 415 416 417 418